Find the critical value for the hypothesis test, given the following. (Give your answer correct to two decimal places.)

Ha: σ1 > σ2, with n1 = 8, n2 = 13, and α = 0.025

F =
Answer #1

n 1 = 8

n 2 = 13

df1 = n 1 - 1 = 7

df2 = n 2 - 1 = 12

\alpha = 0.025

F 7 , 12 , 0.025 = 3.607 = 3.60